OpenStudy (anonymous):

Studies of gymnasts show that their high rate of injuries to the Achilles tendon is due to tensions in the tendon that typically reach 11 times body weight. That force is provided by a pair of muscles, each exerting a force at 20 degrees to the vertical, with their horizontal components opposite. For a 54-kg gymnast, find the force in each of these muscles. Am I starting this right??

OpenStudy (anonymous):

w=m*g \[w = 54kg(9.81m/s^2)\] w = 529.74N  The trouble is, I keep coming up with the wrong numbers and I'm not sure how to proceed...

OpenStudy (anonymous):

Nevermind. Came up with the correct answer and forgot to halve it for the correct answer.

OpenStudy (anonymous):

Rather, came up with the correct *process for* the answer.